OpenStudy (anonymous):

Solve 5+√3x-8=21 A. 88 B. 8 C. 228 D. 7

OpenStudy (shamil98):

Subtract 5 from sides and then square both sides.

OpenStudy (anonymous):

how?

OpenStudy (shamil98):

try doing the work.

OpenStudy (jdoe0001):

\(\bf 5+\sqrt{\square-\square }=\square \qquad \textit{subtract 5 from both}\\ \quad \\ 5+\sqrt{\square-\square }-5=\square -5\qquad \textit{square both sides}\\ \quad \\ (5+\sqrt{\square-\square }-5)^2=(\square -5)^2\)